# FLeNS: Federated Learning with Enhanced Nesterov-Newton Sketch

### v1

@@ I Introduction
Federated Learning (FL) has emerged as an up-and-coming framework for large-scale machine learning tasks, providing substantial benefits in terms of privacy protection and computational efficiency [1, 2, 3, 4, 5]. By enabling decentralized model training without sharing raw data, FL addresses concerns of data sensitivity and ownership. However, a significant challenge in FL is balancing the tradeoff between fast convergence rates and minimizing communication overhead.
@@
In contrast, second-order methods, such as Newton’s method, are known for their superior convergence rates in centralized settings, often achieving linear or super-linear convergence under mild conditions [13, 14]. However, directly applying these methods in FL is impractical due to the excessive communication costs associated with transmitting local Hessian matrices. Various approximations, including BFGS [15], L-BFGS [16], inexact Newton [17], Gauss-Newton [18] and Newton sketch [19], have been proposed to mitigate computational complexity, but adapting them to FL while maintaining communication efficiency remains a challenge.
@@ II Related Work
To accelerate convergence, Newton Sketch [19] was introduced as a second-order method that approximates Newton’s method using sketching techniques to reduce computational and communication burdens. Newton Sketching has been further optimized by methods such as Newton-LESS [21, 22, 23] reducing the effective dimension of the Hessian matrix, reducing communication complexity to $O(kM)$, where $k$ is the sketch size.
@@ III Background
The number of examples on each local machine is denoted by $n_{j}=|D_{j}|$, and the total number of examples is $N=\sum_{j=1}^{m}n_{j}$. The goal is to train a global model using this decentralized data concerning the empirical risk minimization objective.
@@ III-A Centralized Newton’s Method
where $\ell$ is the loss function, $\alpha(w)$ is the regularization term, and $\lambda>0$ is the regularization parameter. Here, $f\in\mathcal{H}_{K}$ belongs to the reproducing kernel Hilbert space (RKHS) defined by a Mercer kernel $K:X\times X\rightarrow\mathbb{R}$. This ensures that the model is defined as:
@@
where $\phi:X\rightarrow\mathcal{H}_{K}$ is an implicit feature mapping.
@@
The time complexity of Newton’s method is $O(NM^{2}t)$, where $M$ is the dimension of the feature space and $t$ is the number of iterations. Newton’s method exhibits a super-linear convergence rate $t=\log(\log(1/\delta))$ for a $\delta$-accurate solution [29], but it incurs high communication and computational costs in federated settings.
@@ V-D Generalization Analysis of FLeNS
Proofs:
  • Lemma 1: Use the strong convexity of $L(w)$ and the boundedness of the gradients. Apply concentration inequalities to bind the difference between the empirical and expected losses.
  • Lemma 2: Utilize McDiarmid’s inequality to bound the deviation of $L(w_{t})$ from its empirical counterpart $L_{S}(w_{t})$. The boundedness of the loss function and its gradients ensures the applicability of the concentration inequality.
@@ VIII Results and Discussions
Figure 1 shows the loss discrepancy (the difference between the loss of the trained model and the optimal solution) over the number of communication rounds for various methods reporting the convergence of the compared methods, and highlighting several key findings. 1) There are significant gaps between the convergence speeds of the proposed method, FLeNS, and existing Newton-type federated learning (FL) methods such as FedNS, FedNew, FedNL, and FedAvg. These gaps validate the accelerated super-linear convergence of FLeNS over other methods. 2) FLeNS converges faster than both FedNS and FedNDES, approaching the performance of FedNewton while requiring less communication and computation. In contrast, FedNew and FedNL converge more slowly, closely resembling the behaviour of FedAvg. 3) Since a lower loss generally correlates with higher predictive accuracy on the test set, we can conclude that FLeNS results in better final predictive accuracy. Despite small sketch sizes, FLeNS demonstrates robustness by maintaining considerable accuracy, making it suitable for scenarios with limited computational resources. 4) FLeNS communicates sketched Hessians of size $k\times k$, denoted as $O(\tilde{d}\lambda^{2})$, significantly reducing per-round communication compared to methods like FedNS and FedNDES. Additionally, Nesterov’s acceleration allows FLeNS to converge faster, meaning it requires fewer iterations to achieve the desired accuracy. By combining lower per-round communication costs with fewer communication rounds, FLeNS results in the lowest total communication cost among the three methods.

### v2

@@ I Introduction
Federated Learning (FL) has emerged as an up-and-coming framework for large-scale machine learning tasks, by providing substantial benefits in terms of privacy protection and computational efficiency [1, 2, 3, 4, 5]. By enabling decentralized model training without sharing raw data, FL addresses concerns of data sensitivity and ownership. However, a significant challenge in FL is balancing the tradeoff between fast convergence rates and minimizing communication overhead.
@@
In contrast, second-order methods, such as Newton’s method, are known for their superior convergence rates in centralized settings, often achieving linear or super-linear convergence under mild conditions [13, 14]. However, directly applying these methods in FL is impractical due to the high communication costs associated with transmitting local Hessian matrices. Various approximations, including BFGS [15], L-BFGS [16], inexact Newton [17], Gauss-Newton [18] and Newton sketch [19], have been proposed to mitigate computational complexity, but adapting them to FL while maintaining communication efficiency remains a challenge.
@@ II Related Work
To accelerate convergence, Newton Sketch [19] was introduced as a second-order method that approximates Newton’s method using sketching techniques to reduce computational and communication burdens. Newton Sketching has been further optimized by methods such as Newton-LESS [21, 22, 23] reducing the effective dimension of the Hessian matrix, reducing communication complexity from $O(M^{2})$ to $O(kM)$, where $k\ll M$ is the sketch size.
@@ III Background
The number of examples on each local machine is denoted by $n_{j}=|D_{j}|$, and the total number of examples is $N=\sum_{j=1}^{m}n_{j}$. The goal is to train a global model using this decentralized data based on the empirical risk minimization objective.
@@ III-A Centralized Newton’s Method
where $\ell$ is the loss function, $\alpha(w)$ is the regularization term, and $\lambda>0$ is the regularization hyparameter. Here, we assume, $f\in\mathcal{H}_{K}$ belongs to the reproducing kernel Hilbert space (RKHS). defined by a Mercer kernel $K:X\times X\rightarrow\mathbb{R}$. This ensures that the model is defined as:
@@
where $\phi:\mathcal{X}\rightarrow\mathcal{H}_{K}$ is the implicit feature mapping.
@@
In federated learning, the system cannot access local training data directly, making it difficult to compute the global gradient $g_{D,t}$ and Hessian matrix $H_{D,t}$ accurately. The total time complexity of Newton’s method is $O(NM^{2}t)$, where $M$ represents the dimension of the feature space and $t$ denotes the number of iterations. Newton’s method achieves a super-linear convergence rate of $t=\log(\log(1/\delta))$ for a $\delta$-accurate solution [29].
@@ VIII Results and Discussions
Figure 1 shows the loss discrepancy (the difference between the loss of the trained model and the optimal solution) over the number of communication rounds for various methods reporting the convergence of the compared methods, and highlighting several key findings. 1) There are significant gaps between the convergence speeds of the proposed method, FLeNS, and existing Newton-type federated learning (FL) methods such as FedNS, FedNDES, FedNew, FedNL, and FedAvg. These gaps validate the accelerated super-linear convergence of FLeNS over other methods. 2) FLeNS converges faster than both FedNS and FedNDES, approaching the performance of FedNewton while requiring less communication and computation. In contrast, FedNew and FedNL converge more slowly, closely resembling the behaviour of FedAvg. 3) Since a lower loss generally correlates with higher predictive accuracy on the test set, we can conclude that FLeNS results in better final predictive accuracy. Despite small sketch sizes, FLeNS demonstrates robustness by maintaining considerable accuracy, making it suitable for scenarios with limited computational resources. 4) FLeNS communicates sketched Hessians of size $k\times k$, denoted as $O(\tilde{d}\lambda^{2})$, significantly reducing per-round communication compared to methods like FedNS and FedNDES. Additionally, Nesterov’s acceleration allows FLeNS to converge faster, meaning it requires fewer iterations to achieve the desired accuracy. By combining lower per-round communication costs with fewer communication rounds, FLeNS results in the lowest total communication cost among the three methods.